# Overcoming the Thermal-Noise Limit of Room-Temperature Microwave Measurements by “Cavity Pre-cooling” with a Low-Noise Amplifier —Application to Time-resolved Electron Paramagnetic Resonance

### v1

@@ Abstract
We demonstrate an inexpensive, bench-top method, which we here name “cavity pre-cooling” (CPC), for removing, just prior to performing a measurement, a large fraction of the deleterious thermal photons that would otherwise occupy the electromagnetic modes of a microwave cavity (or some alternative form of rf resonator) at room temperature. Our method repurposes the input of a commercial HEMT-based low-noise amplifier (LNA) to serve as a photon-absorbing “cold load” that is temporarily over-coupled to the cavity. No isolator is inserted between the LNA and the cavity’s coupling port. In a proof-of-concept experiment, the noise temperature of a monitored microwave mode drops from a little below room temperature to $\sim$108 K. Upon incorporating our pre-coolable cavity into a time-resolved (tr-) EPR spectrometer, a commensurate improvement in the signal-to-noise ratio is observed, corresponding to a factor-of-5 speed up over a conventional tr-EPR measurement at room temperature for the same precision and/or sensitivity. Simulations indicate the feasibility, for realistic Q factors and couplings, of cooling the temperatures of a cavity’s modes down to a few tens of K, and for this coldness to last several tens of $\mu s$ whilst the cavity and its contents are optimally interrogated by a microwave tone or pulse sequence. The method thus provides a generally applicable and extremely convenient approach to improving the sensitivity and/or read-out speed in pulsed or time-resolved EPR spectroscopy, quantum detection and other radiometric measurements performed at or near room temperature. It also provides a first-stage cold reservoir (of microwave photons) for deeper cooling methods to work from.
preprint:
@@ I Introduction
In the face of thermal noise, quantum measurements very often require cryogenic environments to achieve sufficient sensitivity and/or fidelity. In the microwave regime of frequencies, this involves cooling the measuring hardware’s cavities and connecting waveguides down to a temperature of a few tens of mK so as to reliably eliminate the incoherent thermal photons that would otherwise reside within them and wreak havoc. Obeying Bose-Einstein statistics, the expected number of such photons occupying any one of the cavity’s electromagnetic (EM) modes is $\bar{n}=[e^{(hf_{\text{mode}}/k_{\rm B}T})-1]^{-1}$, $f_{\text{mode}}$ being the mode’s frequency. This level of cooling can be achieved, albeit onerously, by installing the hardwave within a deeply cryogenic refrigerator, most often a dilution or adiabatic-demagnetization fridge.
@@
Though currently still a few factors of ten greater in temperature/photon number than what can be achieved through standard mK refrigeration (as involves the removal of thermal phonons from the materials that surround or fill an electromagnetic cavity), recent studies[1, 2, 3, 4, 5] have demonstrated the direct removal of thermal photons from particular microwave modes inside bench-top cavities at room temperature through the stimulated absorption of these photons by an optically spin-polarised medium placed inside the cavity. While obviating the need for physical refrigeration, this hybrid (optical-microwave) approach still requires a sufficiently powerful optical pump source for spin-polarizing the absorbing medium. This results in still quite a bulky and power-hungry system, even when a miniaturized microwave cavity of high magnetic Purcell factor is used to reduce the optical pump power/energy needed[6, 7].
@@
Within microwave instrumentation, low-noise amplifiers (LNAs) are ubiquitously used to boost the power of weak signals. Guided quantitatively by Friis’ formula[8], the deployment of the quietest LNA available at the front end of an instrument’s amplification chain will minimize the amount of deleterious noise added to the measured signal, and so maximize the overall signal-to-noise ratio (SNR) and measurement sensitivity[9]. Compared to their manifest utility in radio-astronomy[10] and deep-space communication[11], the gains in sensitivity from installing a lower-noise (cryogenic) preamplifier into an EPR/NMR spectrometer are found, empirically, to be modest[12, 13, 14], especially when the spectrometer’s microwave/rf cavity is not itself refrigerated, and furthermore dependent on the degree of resonator-amplifier coupling[15].
@@
Various schemes such as Townes’ “Q-multiplier”[16], involving the incorporation of a maser gain medium inside the cavity itself, have been proposed over the years towards improving measurement sensitivity. It was Mollier et al.[17] who cut through the confusion to affirm that, no matter how low-noise the receiving LNA gain medium, the sensitivity of an EPR (or NMR) spectrometer is ultimately limited by thermal noise (i.e. incoherent microwave photons) generated from within the EPR cavity itself. An ultra-low-noise receiving amplifier attached directly to the cavity’s coupling port, or even an ultra-low-noise gain medium resident inside the cavity, will amplify both the signal and this noise equally, thus the SNR won’t improve. As addressed quantitatively by Poole[18] and others[19, 20] since, the only ways to improve sensitivity (for a given sample, completely filling the cavity, at a given temperature) are to either increase the cavity’s quality factor $Q$ (albeit at the expense of measurement bandwidth), so increasing the photon cavity lifetime, or to increase the EPR transition frequency (by applying a strong d.c. magnetic field to the sample), so reducing the number of thermal photons in the cavity’s operational mode. Or else (back to where this argument began:) to physically cool the cavity or otherwise “state-prepare” its operational microwave mode(s).
Figure 1: (a) Conceptual mechanism of cavity pre-cooling (CPC): A microwave resonator is physically maintained at a temperature $T_{\text{hot}}$ (in this article, $T_{\text{hot}}$ equals room temperature, denoted as $T_{\text{0}}$) and supports an electromagnet mode whose photon occupation number is $\bar{n}_{\text{mode}}$. This mode is coupled to a cold load maintained at a lower temperature $T_{\text{cold}}$ with a coupling factor of $\kappa_{\text{over}}$ between the load and the mode. (b) Simplified equivalent electrical circuit: The resonator, drawn as an LCR tank circuit, is over-coupled (right) through a (SPDT) switch to either a cold load (in the form of the input of an LNA) or an open circuit -depending on the position of the switch. The mode’s field amplitude (as generates a voltage across the tank circuit) can be measured by a second LNA through its own coupling port (left, usually set to be at critical coupling with the cooling LNA switched out) whose output is recorded on a digital oscilloscope.
@@
At least for frequencies in the low-GHz region, today’s solid-state LNAs, usually in the form of MMICs or hybrids based on InP (or GaAs) HEMTs or else SiGe HBTs, offer impressively low noise temperatures (referred to input) of 50 K or less -even when operating at room temperature. As mature, COTS technologies, they are readily purchasable. If suitably designed/chosen, the noise temperature of such an LNA will decrease by a further order of magnitude or so upon cooling it to liquid-helium temperature[21, 22]. With such low amplifier noise temperatures available, the sensitivity with which the field-amplitude/energy ($=$ number of photons) inside a room-temperature microwave cavity can be measured is determined almost entirely by the cavity’s own thermal noise (as is associated with the rate at which electromagnetic energy is dissipated inside the cavity) -not the small amount of extra noise imparted by the receiving amplifier chain used to monitor the cavity’s field amplitude (through a port attached to the cavity).
Figure 2: (a) Photograph of the experimental resonator’s insides: a sapphire ring holding a 0.1 % pentacene-doped para-terphenyl (Pc:PTP) crystal, surrounded by the silver-plated walls of a cylindrical brass enclosure. (b) CAD diagram of the same resonator with a two-dimensional axisymmetric simulation of the magnetic flux (white arrows) and magnetic energy density of the TE01δ microwave mode[23] that this resonator supports. (c) Signal-path anatomy of the experimental tr-EPR spectrometer used to demonstrate cavity pre-cooling. This circuit incorporates homodyne detection of the transient EPR signal. The firing of the OPO laser and the closing/opening of the two microwave switches are scheduled via TTL control pulses (generated by two cascaded pulse generators -both a Thandar TG105) as explained in Fig. 5.
@@
As already mentioned above, Wu et al.[1] demonstrated the cooling of an electromagnetic (EM) mode by coupling it to a “spin-cold” reservoir in the form of an optically excited crystal of pentacene-doped para-terphenyl. Ng et al.[2], Blank et al.[24], and Day et al.[5], have since demonstrated quasi-continuous mode cooling using photoexcited negatively charged nitrogen vacancies (NV-) in diamond as a spin-polarised absorber. In a similar vein, Englund, Trusheim and co-workers[3, 4] have exploited mode cooling with NV- diamond to enhance the sensitivity of a magnetometer. All of these works use a cm-sized dielectrically-loaded cavity to support a TE01δ mode affording a high magnetic Purcell factor. The frequency, $f_{0}$, of the this mode is adjusted or else a d.c. magnetic field is applied, so as to align it with the absorptive paramagnetic transition being targeted. Since the linewidth of this transition is narrow (a few MHz for pentacene, a few hundreds of kHz for NV- diamond), these approaches to mode cooling are intrinsically narrow-band.
In microwave radiometry, cooled coaxial or waveguide terminations are used as cryogenic loads[25] to absorb unwanted signals and noise. Such a load in combination with a cryogenic circulator (so forming an isolator) is also often used to protect a sensitive quantum circuit from noise back-flowing from a warmer read-out amplifier. In this context, we point out that the cavity-pre-cooling method introduced here draws parallels with the radiatively cooling of a spin ensemble or a resonator, held physically at liquid-helium temperature, by coupling it electromagnetically to an “absorber” (in the form of either a passive load or active device) at mK temperature[26, 27]. A convenient alternative to a physically cryogenic termination is an active cold load (ACL)[28, 29]) which, though a powered device, mimics the former’s behavior. ACLs can be made from the inputs (gates/bases) of suitably biased low-noise transistors [30, 31]. Though not always as cold (nor as broadband or stable) as true cryogenic terminations, they circumvent the need for refrigeration, enabling the advantageous use of ACLs in mobile applications such as the calibration of earth-monitoring radiometers aboard aircraft or satellites[32].
II Cavity Pre-Cooling
The essential idea of the method introduced here is to temporarily cool the EM modes of a room-temperature cavity by (over-)coupling them, temporarily, to an active cold load in the form of the input of a commercial low-noise amplifier (LNA). The output from this LNA does not need to go anywhere at all; it can simply be sunk into a matched passive load or, if the backwards attenuation of the cooling LNA is insufficient, into a second ACL provided by the input of an additional LNA. The method affords wide-band cooling: it will cool all of the modes lying within the LNA’s bandwidth, provided they are overcoupled to the LNA’s input. The concept is shown in Fig. 1. Here, the input of the cooling LNA is temporarily connected via a low-insertion-loss switch to a large loop of wire inside the cavity, which serves as an over-coupled port. This overcoupling ensures that the temperature of a microwave mode (within the LNA’s input bandwidth) is dominated by the LNA’s input noise temperature as opposed to that associated with the cavity’s own intrinsic dissipation (i.e. the temperature at which energy is dissipated in the metal walls and internal dielectric components of the cavity, namely the cavity’s physical ambient temperature). In our proof-of-principle set up here, the temperature of a particular (“target”) mode is monitored through a second port to which the input of a second LNA is critically coupled (once the cooling LNA on the overcoupled port has been disconnected). The output of this monitoring LNA feeds a homodyne receiver tuned to the target mode’s center frequency.
Figure 3: Simulated average number of thermal photons occupying the cavity’s TE01δ mode and the mode’s corresponding temperature ($T_{\text{mode}}$) as a function of the cold load’s own noise temperature ($T_{\text{cold}}$) and its coupling factor ($\kappa$) to the mode. In our experimental demonstration, this mode’s temperature is reduced from 255.4 K to 108.1 K, the latter temperature corresponding to an average occupancy of 1553 photons.
Details: Our implemented experimental set-up is shown in Fig. 2. However, when solely demonstrating mode cooling, (i) the cavity need not be loaded with an EPR sample, and (ii), the EPR bridge’s circulator can be removed with the monitoring LNA (inside the dashed green box in Fig. 2) connected directly to the critically-coupled port’s stub tuner. Our resonator takes the form of a cylindrical brass cavity, internally plated with silver, whose internal volume is partially filled by a concentrically located ring of monocrystalline sapphire (supplied by J-Crystal Photoelectric Technology, China) whose c-axis lies parallel to the cavity’s axis of rotational symmetry[23]. This ring supports a TE01δ mode at 1.45 GHz, whose intrinsic quality factor $Q_{\text{cav}}$ = 164,000, providing a loaded quality factor at critical coupling ($\kappa_{c}^{0}=1$) of $Q_{\text{crit.}}=Q_{\text{cav}}/2$ = 82,000. The frequency $f_{0}$ and field profile of this mode were accurately modeled using COMSOL Multiphysics[33]. Unlike the superheterodyne receiver (incorporating a narrowly filtered 70-MHz IF strip) used in the previous cavity-cooling works by Wu and Ng et al. [1, 2], we here use a simpler high-gain homodyne receiver to monitor the amplitude/excitation of the TE01δ, recording the receiver’s output into a Tektronix TBS 1102B-EDU oscilloscope.
It is worth pointing out that both the cooling and the monitoring of the TE01δ mode could alternatively be implemented using just a single port attached to a single LNA if this port’s coupling to the mode were made dynamically adjustable, with the LNA alternatively connected (“toggled”), at sufficient speed and glitch-free, between (say) a larger and a smaller coupling loop. However, given that our cavity already had two separate coaxial ports (feedthroughs), it was simpler, with the COTS components immediately available to us, to use both of these ports set to different fixed couplings $\kappa$, namely: (i) an over-coupled “cooling” port ($\kappa_{\text{over}}=3.8$), connected through a low-insertion-loss SPDT switch (viz. a Qorvo RFSW1012) to either the active cold load (viz. the input of a Qorvo QPL9547 LNA mounted on an eval. board) or else an “open” termination -the choice being determined by the control voltage applied to the RFSW1012 switch; and (ii) a “monitoring/interrogation” port, critically coupled ($\kappa_{c}^{0}=1$) to the TE01δ mode through a Maury Microwave 1819A triple-stub tuner, and used to monitor the mode’s field amplitude (and thus noise temperature). These two different ports are depicted, conceptually, in Fig. 1(b); their experimental embodiments are displayed in Figs 2(a) and 2(c). The coupling factors of both ports (cooling and monitoring) as well as the quality factor of the sapphire resonator’s TE01δ mode were accurately determined using a trusty HP8753C vector network analyzer (VNA)[34, 35].
What depth of cooling can one expect? From Fig. 1 (a): The microwave cavity (grey) is kept at an intrinsic physical temperature of $T_{\text{hot}}=T_{\text{0}}$ (orangey brown) but the microwave mode that it supports is also coupled, with a coupling factor of $\kappa_{\text{over}}$, to a cold load (blue), whose effective temperature is $T_{\text{cold}}$. On average, $\bar{n}_{\text{mode}}$ thermal photons will occupy the mode, corresponding to a mode temperature of $T_{\text{mode}}=hf_{\text{mode}}/[k_{\rm B}\textrm{ln}(1+1/\bar{n}_{\text{mode}})]$, where
$T_{\text{mode}}=\frac{T_{\text{0}}+\kappa_{\text{over}}T_{\text{cold}}}{1+\kappa_{\text{over}}}.$ (1)
One immediately sees from this equation that, if the coupling factor $\kappa_{\text{over}}\gg 1$ (corresponding to a substantially overcoupled port), the microwave cavity’s noise temperature will approach that of the cold load, $T_{\text{cold}}$. In practice, the noise temperature on the over-coupled port as seen/felt by a mode will be slightly higher than the cooling LNA’s input noise temperature due to room-temperature insertion losses (I.L.) between the LNA’s input and the cavity. These losses comprise [see Fig. 2(c)] the I.L. of the SPDT switch itself plus those of the two connecting coaxial cables on either side of it. The increased noise temperature generated by them equals $\lambda T_{\text{0}}$, with $\lambda=L_{\text{dB}}/4.34$, $L_{\text{dB}}$ being the overall I.L. in dB[36]. Furthermore, we need to include the (albeit relatively modest) cooling effect of the critically-coupled monitoring LNA (viz. a second, nominally identical, Qorvo QPL9547EVB-01) that is permanently coupled to the mode. Including all of the above additional contributions (as per Eq. S13 of the Supplementary Material), the lowest obtainable mode temperature, $T_{\text{mode}}^{\text{cooled}}$, is predicted to be
$T_{\text{mode}}^{\text{cooled}}=\frac{(1+\kappa_{\text{over}}\lambda)T_{\text{0}}+\kappa_{\text{over}}T_{\text{cold}}+T_{\text{mon}}}{2+\kappa_{\text{over}}},$ (2)
where $T_{\text{mon}}$ is the effective cooling temperature of the monitoring LNA taking into account the unavoidable insertion loss of the triple-stub tuner connecting it to the cavity’s monitoring port, namely $\sim$6 dB, which brings $T_{\text{mon}}$ up to 220.3 K. The expected noise temperature inside the resonator upon pre-cooling, as calculated from Eq. (2) above, is $T_{\text{mode}}^{\text{cooled}}$ = 108.1 K.
Figure 4: Increase in measured noise over time as the experimental cavity’s monitored microwave mode warms up from “pre-cooled” to ambient conditions. The two black dashed lines track the envelope of the voltage noise. Upon curve-fitting to this envelope (see Fig. 15 in the Supplementary Material), the characteristic time of the warming up is determined to be $\tau_{\text{warm up}}^{\text{exp.}}\approx 9\pm 1\mu$s.
Figure 5: Cavity pre-cooling as applied to time-resolved EPR: In Step 1, an LNA connected to an over-coupled cooling port extracts thermal noise (energy) from the microwave cavity (represented here as an LCR circuit); this also pulls the cavity mode’s resonance frequency $f_{0}$ –slightly. Next (Step 2), a switch disconnects the cooling LNA, bringing the mode back on resonance (i.e., its frequency coincident with the EPR transition’s line center at 1.4495 GHz). Instantaneously with this microwave switching, a q-switched optical pulse (around 7 nanoseconds in duration) of 532-nm pump light is shone on the EPR sample within the the cavity. Subsequently (Step 3), an interrogating coherent microwave tone is applied via a circulator to the cavity’s monitoring port, as shown in Fig. 2 (c), which, upon re-emerging from the cavity through the circulator’s third arm, provides a tr-EPR signal characterizing the EPR sample.
It is worth pointing out how the monitoring LNA itself helps to cool the mode that it monitors. This would not happen if, as is often done, a microwave isolator were inserted between the monitoring LNA and the cavity -so as to “protect” the cavity from external noise. Noise (and thus sensitivity) wise, it is always advantageous to connect the monitoring LNA directly to the cavity so long as its input noise temperature is lower that the physical temperature of the isolator’s internal termination.
Figure 6: (a) Zero-field tr-EPR signal of Pc:PTP’s $T_{x}-T_{z}$ transition without (red) and with (blue) pre-cooling. (b) Thermal noise (in the time domain) extracted from each of these ZF tr-EPR signals and, below, their corresponding spectral densities; as before, blue and red correspond to the pre-cooled (with) and ambient (without) conditions, respectively. (c) Simplified Jablonski (= energy-level) diagram of the ZF sublevels ($T_{x}$, $T_{y}$, $T_{z}$) of pentacene’s triplet ground state, where D and E denote the ZF splitting (ZFS) parameters. (d) Simulated noise power reduction ($\Delta P$) at the homodyne receiver’s input as a function of the noise temperature of the EPR cavity’s operational mode, $T_{\text{mode}}$. Since its uncertainty is lower, the estimated depth of cooling presented here ($\Delta P_{\text{exp-Cool}}$) corresponds to that achieved by CPC alone (see Table S2 in the Supplementary Material). The depth’s standard error is indicated by dashed cornflower-blue lines.
The reduction in thermal noise will persist so long as the cooling LNA is connected. Upon disconnecting it (so connecting the over-coupling loop to a reflective “open”), the mode regains its higher quality factor, namely $Q_{\text{crit.}}$, and warms up on a timescale set by the mode’s photon lifetime[37], namely $\tau_{\text{warm up}}=\frac{Q_{\text{crit.}}}{2\pi f_{0}}.$ Since the monitoring port is critically coupled to the TE01δ mode, this mode’s temperature will settle at the (equally weighted) average of $T_{\text{hot}}=290$ K and $T_{\text{mon}}$, namely $T_{\text{mode}}^{\text{ambient}}=255.4$ K for our particular experiment.
III Results
Preliminaries: We determined the noise characteristics of our cooling LNA, namely a Qorvo QPL9547EVB-01 operating at room temperature, using the industry-standard Y-factor method[38]. At 1.45 GHz, it exhibits an impressive noise figure of just 0.26 dB (referenced to 290 K), equivalent to a noise temperature of 18.2 K; see Table S1 in the Supplementary Material for this LNA’s full noise parameters. Experimentally, we observed a reproducible transient on the EPR (homodyne) signal as the SPDT switch opened and disconnected the pre-cooling LNA from the microwave cavity. This artefact (lasting less than 3 $\mu$s in our experiment) could be substantially removed by measuring it for several switch openings then subtracting the average from our measured traces.
Mode Cooling: The effect of connecting then disconnecting the over-coupled cooling LNA to our cavity’s TE01δ mode at $f_{0}=$ 1.45 GHz is shown in Fig. 4. From the difference in the time average of the measured voltage squared for the mode under cooled and ambient conditions, the depth of cooling is estimated to be $\Delta P_{\text{exp-Cool}}=-3.5\pm 0.4$ dB. This agrees with both our theoretical estimate $\Delta P_{\text{sim}}$ and our analysis using the noise spectrum from experimental tr-EPR data, namely $\Delta P_{\text{exp-EPR}}$, as shown in Fig. 6(b). All our estimates are summarized in Table T2 of the Supplementary Material. Upon curve-fitting to the rise in the measured voltage squared as a function of time, the characteristic time quantifying the persistence of the mode’s coldness after disconnection from the cooling LNA, is estimated to be $\tau_{\text{warm up}}^{\text{exp.}}=9\pm 1~{}\mu$s; see Fig. 15 in the Supplementary Material.
Applicaton to EPR: The protocol used to implement cavity pre-cooling (CPC) is depicted in Fig. 5 and consists of three stages: initial noise extraction via an over-coupled port connected to an active cold load (in the form of the input of an LNA), followed by the disconnection of the ACL, re-tuning the operational mode of the cavity back onto the line center of the target EPR transition, combined with simultaneous laser and microwave excitation to elicit the tr-EPR signal. The process culminates with the signal’s capture through homodyne detection.

### v2

# Overcoming the Thermal-Noise Limit of Microwave Measurements by Pre-cooling with an Active Cold Load

@@ Abstract
We introduce a method, which we here name “active pre-cooling” (APC), for removing, just prior to performing a measurement, a large fraction of the deleterious thermal photons that would otherwise occupy the electromagnetic modes of a microwave cavity or some alternative form of radio-frequency resonator. The removal is achieved by temporarily over-coupling the cavity’s modes to an active cold load (ACL). We report a room-temperature bench-top demonstration of the method, where this load takes the form of the input of a commercial low-noise amplifier (LNA). No isolator is inserted between the LNA’s input and the cavity’s coupling port. The noise temperature of a monitored microwave mode drops to $\sim$123 K. Upon incorporating our pre-coolable cavity into a time-resolved (tr-) EPR spectrometer, a commensurate improvement in the signal-to-noise ratio is observed, corresponding to a factor-of-5 speed up over a conventional tr-EPR measurement at room temperature for the same precision and/or sensitivity. Modeling indicates the feasibility, for realistic mode quality factors and couplings, of cooling the room-temperature cavity’s modes down to a few tens of K, and for this coldness to last several tens of $\mu s$, whilst the cavity and its contents are optimally interrogated by a microwave tone or pulse sequence. The method thus provides a simple and generally applicable approach to improving the sensitivity and/or read-out speed in pulsed and time-resolved EPR spectroscopy, quantum detection and other radiometric measurements. It provides a first-stage cold reservoir (of microwave photons) for other, deeper cooling methods to work from and, through the use of cryogenic ACLs (realized, in the first instance, as the inputs of existing cryogenic low-noise amplifiers), the method could itself be directly extended to lower temperature regimes.
@@ I Introduction
In the face of thermal noise, quantum measurements very often require cryogenic environments to achieve sufficient sensitivity and/or fidelity. In the microwave regime of frequencies, this involves cooling the hardware’s cavities and connecting waveguides down to a temperature of a few tens of millikelvin (mK) so as to reliably eliminate the incoherent thermal photons that would otherwise reside within them and cause havoc. Obeying Bose-Einstein statistics, the expected number of such photons occupying any one of the cavity’s electromagnetic (EM) modes is given by $\bar{n}=[e^{(hf_{\text{mode}}/k_{\rm B}T})-1]^{-1}$, where $f_{\text{mode}}$ is the mode’s frequency. This level of cooling can be achieved, albeit onerously, by installing the hardware within a deeply cryogenic refrigerator, most often a dilution or adiabatic-demagnetization fridge.
@@
Though currently still a few factors of ten greater in temperature/photon number than what can be achieved through standard mK refrigeration (as involves the removal of thermal phonons from the materials that surround or fill an electromagnetic cavity), recent studiesWu et al. (2021); Ng et al. (2021); Fahey et al. (2023); Wang et al. (2024); Day et al. (2024) have demonstrated the direct removal of thermal photons from particular microwave modes inside bench-top cavities at room temperature through the stimulated absorption of these photons by an optically spin-polarised medium placed inside the cavity. While obviating the need for physical refrigeration, this hybrid (optical-microwave) approach still requires a sufficiently powerful optical pump source for spin-polarizing the absorbing medium. This results in still quite a bulky and power-hungry system, even when a miniaturized microwave cavity of high magnetic Purcell factorBreeze et al. (2015); Fleury et al. (2023) is used to reduce the optical pump power/energy needed.
@@
Within microwave instrumentation, electrically powered low-noise amplifiers (LNAs) are ubiquitously used to increase the amplitude (and thus power) of weak signals. Guided quantitatively by Friis’ formulaFriis (1944), the deployment of the quietest LNA available at the front end of an instrument’s amplification chain will minimize the amount of deleterious noise added to the measured signal, thereby maximize the overall signal-to-noise ratio (SNR) and measurement sensitivityŠimėnas et al. (2021). Compared to their manifest utility in radio-astronomyGiordmaine et al. (1959) and deep-space communicationClauss and Shell (2008), the gains in sensitivity from installing a lower-noise (cryogenic) preamplifier into an electron paramagnetic resonance (EPR) or nuclear magnetic resonance (NMR) spectrometer are found, empirically, to be both (i) modestCollier et al. (1968); Narkowicz et al. (2013); Kalendra et al. (2023), especially when the spectrometer’s microwave or radio-frequency (RF) cavity is not itself refrigerated, and (ii) dependent on the degree of cavity-amplifier couplingPfenninger et al. (1995).
@@
Various schemes such as Townes’ “Q-multiplier”Townes (1960), involving the incorporation of a maser gain medium inside the cavity itself, have been proposed over the years towards improving measurement sensitivity. It was Mollier et al.Mollier et al. (1973) who cut through the confusion to affirm that, no matter how low-noise the receiving LNA gain medium, the sensitivity of an EPR (or NMR) spectrometer is ultimately limited by thermal noise (i.e. incoherent microwave photons) generated from within the EPR cavity itself. An ultra-low-noise receiving amplifier attached directly to the cavity’s coupling port, or even an ultra-low-noise gain medium resident inside the cavity, will amplify both the signal and this noise equally; thus, the SNR won’t improve. As addressed quantitatively by PoolePoole and Farach (2019) and othersHill and Richards (1968); Rinard et al. (1999) since, the only ways to improve sensitivity (for a given sample, completely filling the cavity, at a given temperature) are to either increase the cavity’s quality factor $Q$ (albeit at the expense of measurement bandwidth), so increasing the photon cavity lifetime, or to increase the EPR transition frequency (by applying a strong d.c. magnetic field to the sample), so reducing the number of thermal photons in the cavity’s operational mode. Or else (back to where this argument began:) to physically cool the cavity or otherwise “state-prepare” its operational microwave mode(s).
@@
At least for frequencies in the low-GHz region, today’s solid-state LNAs, in the form of monolithic microwave integrated circuits (MMICs) or else as hybrids composed of interconnected discrete components, containing either InP- or GaAs-based high-electron-mobility transistors (HEMTs) or else SiGe-based heterojunction bipolar transistors (HBTs), offer impressively low noise temperatures (referred to input) of 50 K or less –even when operating at room temperature. As mature commercial-off-the-shelf (COTS) components, they are readily purchasable. If suitably designed/chosen, the noise temperature of such an LNA will decrease by a further order of magnitude or so upon cooling to liquid-helium temperatureWeinreb et al. (1982); Pospieszalski (2024). With such low amplifier noise temperatures available, the sensitivity with which the field-amplitude/energy ($=$ number of photons) inside a room-temperature microwave cavity can be measured is determined almost entirely by the cavity’s own thermal noise (as is associated with the rate at which electromagnetic energy is dissipated inside the cavity) –not the small amount of extra noise imparted by the receiving amplifier chain used to monitor the cavity’s field amplitude through a port attached to the cavity.
@@
As already mentioned above, Wu et al.Wu et al. (2021) demonstrated the cooling of an electromagnetic (EM) mode by coupling it to a “spin-cold” reservoir in the form of an optically excited crystal of pentacene-doped para-terphenyl (Pc:PTP). Ng et al.Ng et al. (2021), Blank et al.Blank et al. (2023), and Day et al.Day et al. (2024), have since demonstrated quasi-continuous mode cooling using photoexcited negatively charged nitrogen vacancies (NV-) in diamond as a spin-polarised absorber. In a similar vein, Englund, Trusheim and co-workersFahey et al. (2023); Wang et al. (2024) have exploited mode cooling with NV- diamond to enhance the sensitivity of a magnetometer. All of these works use a cm-sized “3-D” dielectrically-loaded cavity to support a TE01δ mode affording a high magnetic Purcell factorBreeze et al. (2015). The frequency $f_{0}$ of this mode is adjusted, or a d.c. magnetic field is applied, to align it with the absorptive paramagnetic transition being targeted. Since the linewidth of this transition is narrow (a few MHz for pentacene, a few hundreds of kHz for NV- diamond), these approaches to mode cooling are intrinsically narrow-band.
II Active Cold Loads
In microwave radiometry, cooled coaxial or waveguide terminations, also known as “cold loads”, are used both as sources of noise and as absorbersGervasi et al. (1995) of unwanted signals. Such a cryogenic load is a passive component requiring no source of power. To be useful as an accurately quantified noise source or “reference”, it must be anchored thermally to a heat bath (such as a “cold plate” within a cryostat) whose temperature is accurately known. A cryogenic load can be connected to one of the ports of a (cryogenic) circulator to form an isolator. At least within the circulator’s operational bandwidth, this isolator can be used to protect a sensitive quantum circuit from backwardly propagating thermal noise and/or coherent signals unintentionally generated by reflections further up/along the signal path. In this context, we point out that the pre-cooling method introduced here draws parallels with the radiative cooling of a spin ensemble or resonator, held physically at liquid-helium temperature, by coupling it electromagnetically to an absorber (in the form of either a passive load or active device) held at mK temperatureAlbanese et al. (2020); Wang et al. (2021).
Figure 1: Paradigm of a cavity that is externally coupled to an active cold load (ACL) in the form of the input of a low-noise amplifier (LNA) whose output in terminated.
A convenient alternative to a physically cryogenic termination is an active cryogenic load (ACL)Frater and Williams (1981); Forward and Terry (2010), which, though a powered device, mimics the former’s behavior. ACLs can be made from the inputs (gates/bases) of suitably biased low-noise transistors configured as single-port terminated amplifiersDunleavy et al. (1997); Leynia de la Jarrige et al. (2010); Weissbrodt (2017). Though not always cold enough (nor broad-band or stable enough) to replace truly cryogenic terminations in all applications, they can be used advantageously in airborne and space applications to calibrate (or at least drift-correct) earth-monitoring radiometersShou et al. (2015); Weissbrodt (2017), dispensing with the encumbrance of mobile refrigeration and/or the operational need to “roll” the instrument’s platform to point the radiometer at dark sky (i.e., at 2.7 K cosmic background radiation –serving as an absolute temperature reference).
Acknowledging that other approaches existClerk et al. (2010), we here apply the so-called “wave approach” to analyse the generation and flow of noise in realistic microwave circuits (see chapter 8 of Ref. Siegman (1964), or Refs. Meys (1978); Dorđević et al. (2015)). Through the construction of flow graphs combined with the application of Mason’s “non-touching loop” ruleHewlett Packard / Agilent (1990); Hecken (1981), this diagrammatic approach readily encompasses the frequently encountered complexity of what a laser physicist would describe as “etalon effects”, where simultaneously forward- and back-propagating noise waves form (partial) standing waves along sections of transmission line between two or more impedance mismatches (the equivalent of partially reflecting mirrors functioning as beamsplitters/combiners). The available noise temperature from the input of a low-noise amplifier (of sufficiently high gain) that is terminated by a matched and sufficiently cold load can thereby be quantified asMeys (1978); Escotte (2015):
$T_{\text{cold}}=-T_{\text{min}}+4T_{0}\frac{R_{\text{n}}}{Z_{0}}\frac{|1-S_{11}\Gamma_{\text{opt}}|^{2}}{(1-|S_{11}|^{2})|1+\Gamma_{\text{opt}}|^{2}},$ (1)